40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Wed Sep 28 04:27:33 2016, ericq, Update, General, mucking about alscheck.pdfweird_olpit.pdf
    Reply  Thu Sep 29 09:49:53 2016, rana, Update, General, mucking about 
    Reply  Sat Oct 1 10:03:28 2016, ericq, Update, General, mucking about 
       Reply  Tue Oct 4 02:59:48 2016, ericq, Update, General, mucking about 2016-10-04_ALSXspectra.pdf
Message ID: 12529     Entry time: Tue Oct 4 02:59:48 2016     In reply to: 12527
Author: ericq 
Type: Update 
Category: General 
Subject: mucking about 

[ericq, gautam]

We poked around trying to figure out the X PDH situation. In brief, the glitchiness comes and goes, not sure what causes it. Tried temp servo on/off and flow bench fan on/off. Gautam placed a PD to pick off the pre-doubler AUX X IR light to see if there is some intermittant intensity fluctuation overnight. During non-glitchy times, ALSX noise profile doesn't look too crazy, but some new peak around 80Hz and somewhat elevated noise compared to historical levels above 100Hz. It's all coherent with the PDH control up there though, and still looks like smooth frequency noise...

NB: The IR intensity monitoring PD is temporarily using the high gain Transmon PD ADC channel, and is thus the source of the signal at C1:LSC-TRY_OUT_DQ. If you want to IR lock the X arm, you must change the transmon PD triggering to use the QPD.

Attachment 1: 2016-10-04_ALSXspectra.pdf  75 kB  | Hide | Hide all
2016-10-04_ALSXspectra.pdf
ELOG V3.1.3-